# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ 3 Estimation and Inference
We now move to estimation of the quantities defined in the previous section using data from unit-level randomized experiments. We first derive the the limiting distribution of a differences-in-means estimator when data is generated from a standard RCT that is run on the entire market. This estimator is consistent for the average direct effect, but its asymptotic variance depends on price-elasticity terms that cannot be estimated using data from an RCT that randomizes treatment only. We show that if confidence intervals are constructed using a variance estimator based on the asymptotic variance without price interference, then coverage for the population direct effect will not be asymptotically exact.
In order to perform inference for the direct effect and to perform estimation and inference for the indirect effect, we introduce an augmented randomized trial, with Bernoulli randomization for the treatment and local perturbation of market prices to estimate price elasticities. The estimator for the direct effect remains unchanged from the standard RCT. Algorithmically, the estimator for the indirect effect looks algorithmically like a combination of differences-in-means estimator and an instrumental variables estimator. Under the augmented randomized experiment, we show how to construct asymptotically valid confidence intervals for population versions of both effects. In this section, we assume that experiments are run on the entire market of $n$ participants. However, as described in more detail in Appendix D.3, the extension to settings where only a sub-sample of the market is observed is straightforward.
@@ 3.1 Estimation for the Direct Effect in a Standard RCT
Our first result is that—even under interference that occurs through market prices—the differences-in-means estimator converges to the population direct effect at a $\sqrt{n}$ rate and is asymptotically normal. The variance of the difference-in-means estimator depends both on the variance of the potential outcomes—as in a standard no-interference setting—as well as an additional term ($\Delta_{i}(w,p^{*}_{\pi})$) that is a function of net demand and price sensitivity of outcomes and net demand. [^3] Our finding that the difference in means is consistent for the ADE mirrors general findings in Sävje et al. (2021); however, the fact that we get a $\sqrt{n}$ rate of convergence and a central limit theorem depends on our marketplace model.
@@
Under the conditions of Theorem 2 and with uniform randomization probabilities $\pi_{i}=\pi$ for all $i$, when treatment is allocated according to Design 1,
@@
Although we have shown that the familiar difference-in-means is asymptotically normal around the ADE, the above result also implies that the usual asymptotic variance derived under the no-interference setting does not match $\sigma^{2}_{D}$ in our market equilibrium model. Instead, to construct confidence intervals that are asymptotically exact for the population direct effect, we need estimates of the price elasticities that appear in the $\Delta_{i}(w,p^{*}_{\pi})$ terms. In the next section, we will develop an augmented randomized experiment that allows us to estimate these price sensitivities as well as the AIE, which is not estimable with treatment randomization only.
Finally, we note that Theorem 5 quantifies the errors of $\hat{\tau}_{\text{ADE}}$ as an estimator of the population estimand $\tau^{*}_{\text{ADE}}$. In some settings, however, it may also be of interest to provide inference about the sample direct effect $\tau_{\text{ADE}}$. Combining Theorems 3 and 5, we can verify that
@@
The asymptotic variance here still depends on the market equilibrium effects; furthermore, it depends on the correlation of $\varepsilon_{i}(1)$ and $\varepsilon_{i}(0)$ and so is generally not identified. However, the following result shows that, in an extension of the classic result of Neyman (1923), $\sigma_{D}$ is a conservative upper bound for $\bar{\sigma}_{D}$ and so confidence intervals built using Theorem 5 that are exact for $\tau^{*}_{\text{ADE}}$ will also be conservative for $\tau_{\text{ADE}}$.
@@ 3.2 Augmented Randomized Experiment
In order to estimate spillover effects or the build confidence intervals for the average direct effect, we need to estimate price elasticities. Here, we do so by considering an augmented experiment that introduces small, unit level price perturbation into our randomization design. The net demand and outcome processes remain the same as in the previous section, but prices are shifted at an individual level by $U_{i}$ for individual $i$.
Design 2.
Augmented Randomized Trial. Treatment is assigned based on a uniform randomization probability, so $W_{i}\sim\text{Bernoulli}(\pi)$, where there exists $\eta>0$ such that $\eta\leq\pi\leq 1-\eta$. For $i\in\{1,\,\ldots,n\}$, $j\in\{1,\,\ldots,J\}$, $U_{ij}$ is drawn independently and uniformly at random from $\{-h_{n},+h_{n}\}$, where $h_{n}=cn^{-\alpha}$ with $\frac{1}{4}<\alpha<\frac{1}{2}$ and $c$ is a finite constant.
@@
The price perturbations introduced by this experiment can be interpreted as instruments for estimating price derivatives locally. They shift demand and supply in the neighborhood of the market-clearing price, and since they are randomized, they are independent of other variables that determine demand or outcomes. In a typical two-sided market, implementing the experiment requires introducing a small random discount, fee, or subsidy to both producers and consumers for any products where the treatment is expected to significantly affect supply or demand. Our theory requires that the size of the price perturbation decreases with sample size. For a fixed market size, as the perturbation size increases, the variance of $\hat{\tau}_{\text{AIE}}$ decreases, but large price perturbations distort production and would be likely be viewed unfavorably by various market stakeholders. The simplest guidance for choosing the size of the price perturbation is to choose a value of $c$ so that for the sample size of the experiment, the size of the price perturbation is noticeable to market participants but limited to a small percentage of the current market price (e.g. 1-2%). For markets where it is not possible for reputational or legal reasons to introduce any individual-level price variation, the components of $\tau^{*}_{\text{AIE}}$ can be estimated using price variation across products or time instead, at the cost of introducing additional structural assumptions on the environment.
The experiment affects the equilibrium in finite samples, since the price perturbations affect the market clearing condition in Assumption 1. However, the asymptotically linear expansion in Theorem 1 holds under Design 1 and Design 2, so the limiting price is the same under both designs. In a qualitative sense, this result means that the price perturbations are unlikely to meaningfully affect healthy operation of the marketplace. [^4] The fact that we only consider small, unobstrusive perturbations to the market also means we do not get to observe behaviors at prices far from the large-sample equilibrium $p^{*}_{\pi}$, and so cannot identify or consistently estimate full demand curves.
We now develop a point estimator for the indirect effect. Our proposal consists of a combination of difference-in-means estimators and estimators of price elasticities. The strategy is plug-in estimation based on the functional form for $\tau^{*}_{\text{AIE}}$ derived in Theorem 2. Let $\bm{Y}$ be the $n$-length vector of outcomes, $\bm{U}$ is the $n\times J$ matrix of price perturbations, and $\bm{Z}$ is the $n\times J$ matrix of net demand observations. The estimator is
@@
Our next result provides a central limit theorem for our indirect effect estimator. This estimator converges at a rate that depends on the magnitude of the price perturbations $h_{n}$, and is always slower than the $\sqrt{n}$-rate obtained for the direct effect. In deriving this result, it is helpful to note that algorithmically $\hat{\tau}_{\text{AIE}}$ is the product of a rescaled instrumental variables estimator and a differences in mean estimator. We can then use standard results on the asymptotic behavior of IV estimators to guide our analysis.
@@
Suppose the Assumptions of Theorem 2 hold and that treatment is assigned according to Design 2. Then, the estimated average indirect effect can be written as:
@@ 4.1 Definitions of Conditional Treatment Effects Under Interference
is the expected effect of treating an individual with covariate value $x$ on everyone else’s outcomes in a system of $n$ individuals. These estimands are non-random quantities that depend both on the population distribution and on the sample size. They are policy-relevant for settings where we expect to deploy a policy in a market of similar size and composition to the observed market. When interference is removed, the CADE is equal to the CATE, and the CAIE is zero.
@@
The next step, in Theorem 10, is to derive the population versions of the CADE and CAIE. These are purely population estimands, defined as the limit of $\bar{\tau}_{\text{CADE}}$ and $\bar{\tau}_{\text{CAIE}}$ as the market size grows to infinity. These estimands are a relevant policy counterfactual if we are interested in deploying a policy to a market that is much larger, but otherwise similar to the market being observed. The population estimands are also useful for deriving estimators that are consistent for both the large-market and finite-market estimands. Theorem 10 shows that the population estimands have a simple and interpretable form and implies that estimators for one class of estimands are consistent for the other.
@@
The population estimands have a simple form that suggest estimation strategies for the estimands $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$. The limit of the CADE is the direct treatment effect on outcomes conditional on $x$, holding the equilibrium price fixed. The limit of the CAIE is the direct treatment effect on net demand conditional on $x$, multiplied by an elasticity correction that does not depend on $x$. The augmented randomized experiment from Design 2 can be used to estimate the elasticity corrections and the conditional average treatment effects required to estimate $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$.
@@ 4.2 Equilibrium-Stable Targeting
The above reveals that we can solve for the optimal price-stable policy without access to price elasticities. Thus, we can obtain a plug-in estimate for the optimal rule using CADE estimates derived from an RCT without price perturbations as in Design 1; recall that the CADE itself can be identified without price perturbations (Theorem 11).
The characterization in Proposition 12 involves a functional linear program. We show below, however, that its solution takes on a simple parametric form: The optimal price-stable policy is a thresholding rule that compares the CADE for the outcomes against a shadow-cost of net demand effects $c\cdot\tau_{\text{CADE}}^{*,z}(x)$, where $c\in\mathbb{R}^{J}$ can be interpreted as a shadow price vector.
@@
Under the conditions of Proposition 12, the asymptotic equilibrium-stable targeting problem (36) admits an optimal solution of the form
$\nu_{c,a}(x)=\begin{cases}1&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)>c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ a&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)=c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ 0&\text{else,}\end{cases}$ (37)
for some $c\in\mathbb{R^{J}}$ and $a\in[0,1]$.
@@ 5 Example: Cash-Transfer Experiments
The authors argue that this is phenomenon is due to spillover effects, which largely arise from equilibrium effects in the market for perishable protein sources. Specifically, the authors provide evidence that eggs, which are a commonly consumed perishable protein source in the Philippines, have higher prices in treated compared to control villages where saturation of eligible households is high. This price effect is explained by the increase in demand for protein from eligible families and inelastic supply in remote villages. The higher prices in treated villages lead ineligible children to consume less protein, which explains a significant proportion of the negative effect of the cash transfer program on ineligible children. It thus appears that Assumption 1 is reasonable, in that the spillovers described in detail by the authors are restricted to those that occur through market prices.
@@ 5.1 Treatment Effects in a Village Economy
Under the assumption that markets in remote villages are isolated from one another, the village-level randomized experiment allowed Filmer et al. (2023) to recover treatment effects that include equilibrium effects. The total effect $\tau_{\text{TOT}}$ could be estimated using data from the cluster-randomized experiment by comparing the average outcomes of all children in treated compared to control villages. In this paper, in contrast, our focus is on randomized experiments where only a single market is observed, but randomization is possible at the household level. While our econometric setting is different from that of Filmer et al. (2023), we can use their dataset to conduct a calibrated evaluation of our approach. We first use moments from the cluster-randomized experiment Filmer et al. (2023) to estimate a model of supply, demand, and children’s outcomes in a single remote village market. Then, we use this model to simulate a single-village, unit-randomized experiment, and assess the ability of our estimators of the average direct and indirect treatment effects to recover equilibrium effects from this simulated experiment.
@@
Given a treatment vector $\bm{W}$, the realized household-level outcomes are $Y_{h}(\bm{W})=Y_{h}(W_{h},P(\bm{W}))$ and household level net demand is $Z_{h}(\bm{W})=D_{h}(\bm{W},P(\bm{W}))-\frac{1}{n_{h}}S(P(\bm{W}))$.
The 10 parameters in this model are identified by 10 moments from a subset of the data in Filmer et al. (2023), restricted to households in villages above the 70th percentile of the remoteness index defined in the paper, where equilibrium effects were most meaningful. The estimated model matches these 10 moments:
@@
In Table 1, we show that both the direct and the indirect effect are estimable in finite samples using our proposed unit level augmented experiment. We report the results of a Monte Carlo simulation with 1,000 repetitions and a sample size of 2,000 households to evaluate the bias, variance, and coverage properties of the estimators and confidence intervals for $\hat{\tau}_{\text{ADE}}$ and $\hat{\tau}_{\text{AIE}}$. The table illustrates that the bias of the average direct effect and indirect effect estimator based on the augmented randomize experiment are low in finite samples. Furthermore, the coverage when $\hat{\sigma}_{D}$ and $\hat{\sigma}_{I}$ are used to construct confidence intervals in finite samples is slightly above the asymptotic confidence level of 95% for both population estimands. As expected from Theorem 3, intervals that target $\tau^{*}_{\text{ADE}}$ are conservative for $\tau_{\text{ADE}}$. For the AIE, in simulations the coverage for the sample estimand is also more conservative than for the population estimand.

### v4

@@ 3 Estimation and Inference
We now move to estimation of the quantities defined in the previous section using data from unit-level randomized experiments. We first derive the limiting distribution of a differences-in-means estimator that is valid when data is generated from a standard RCT that is run on an entire market. This estimator is consistent for the average direct effect, but its asymptotic variance depends on price-elasticity terms that cannot be estimated using data from an RCT that randomizes treatment only. We show that if confidence intervals are constructed using a variance estimator based on the asymptotic variance without price interference, then coverage for the population direct effect will not be asymptotically exact.
In order to perform inference for the direct effect and to perform estimation and inference for the indirect effect, we need non-zero price shifters. Algorithmically, the estimator for the indirect effect looks like a combination of differences-in-means estimator and an instrumental variables estimator. Under an augmented randomized experiment, we show how to construct asymptotically valid confidence intervals for population versions of both effects. In this section, we assume that experiments are run on the entire market of $n$ participants. However, as described in more detail in Appendix D.4, the extension to settings where only a sub-sample of the market is observed is straightforward.
@@ 3.1 Estimation for the Direct Effect in a Standard RCT
Our first result is that—even under interference that occurs through market prices—the standard differences-in-means estimator converges to the population direct effect at a $\sqrt{n}$ rate and is asymptotically normal. The variance of the difference-in-means estimator depends both on the variance of the potential outcomes—as in a standard no-interference setting—as well as an additional term ($\Delta_{i}(w,p^{*}_{\pi})$) that is a function of net demand and price sensitivity of outcomes and net demand.  [^6] Our finding that the difference in means is consistent for the ADE mirrors general findings in Sävje et al. (2021); however, the fact that we get a $\sqrt{n}$ rate of convergence and a central limit theorem depends on our marketplace model.
@@
Under the conditions of Theorem 2 and with uniform treatment randomization probabilities $\pi_{i}=\pi$ for all $i$, under Design 1 or Design 2,
@@
Although we have shown that the familiar difference-in-means is asymptotically normal around the ADE, the above result also implies that the usual asymptotic variance derived under the no-interference setting does not match $\sigma^{2}_{D}$ in our market equilibrium model, unless $\Delta_{i}(W_{i},p^{*}_{\pi})=0$. This term is zero if there is homogeneity in price derivatives under treatment and control, so that $\nabla_{p}[y(1,p)-y(0,p)]=0$. In general, to construct confidence intervals that are asymptotically exact for the population direct effect, we need estimates of the price elasticities that appear in the $\Delta_{i}(w,p^{*}_{\pi})$ terms. In the next section, we will show how non-zero price perturbations under Design 2 allows us to estimate these price sensitivities as well as the AIE, which is not estimable with treatment randomization only.
Finally, we note that Theorem 5 quantifies the errors of $\hat{\tau}_{\text{ADE}}$ as an estimator of the population estimand $\tau^{*}_{\text{ADE}}$. In many settings, however, it is of primary interest to provide inference about the sample direct effect $\tau_{\text{ADE}}$, which is defined in a finite-sized market. Combining Theorems 3 and 5, we can verify that
@@
The asymptotic variance here still depends on the market equilibrium effects; furthermore, it depends on the correlation of $\varepsilon_{i}(1)$ and $\varepsilon_{i}(0)$ and so is generally not identified. However, the following result shows that, in an extension of the classic result of Neyman (1923), $\sigma_{D}$ is a conservative upper bound for $\bar{\sigma}_{D}$ and so confidence intervals built using Theorem 5 that are exact for $\tau^{*}_{\text{ADE}}$ will also be conservative for $\tau_{\text{ADE}}$. [^7] If $\tau_{\text{ADE}}$ is the only estimand of interest, we can adapt results from Aronow et al. (2014) to derive a tighter bound for $\bar{\sigma}^{2}_{D}$ that is estimable using price perturbations as discussed in the following section; see Appendix D.2 for details.
@@ 3.2 Augmented Randomized Experiment
We now consider use of randomized individual-level price-shifters as defined in Design 2 to accomplish further estimation and inferential tasks. These price perturbations can be interpreted as experimenter-created instruments for estimating price derivatives locally. They shift demand and supply in the neighborhood of the market-clearing price, and since they are randomized, they are independent of other variables that determine demand or outcomes. [^8] The fact that we only consider small, unobstrusive perturbations to the market also means we do not get to observe behaviors at prices far from the large-sample equilibrium $p^{*}_{\pi}$, and so cannot identify or consistently estimate full demand curves. In a typical two-sided market, implementing the experiment requires introducing a small random discount, fee, or subsidy to both producers and consumers for any products where the treatment is expected to significantly affect supply or demand. [^9] Our theory requires that the size of the price perturbation decreases with sample size. For a fixed market size, as the perturbation size increases, the variance of $\hat{\tau}_{\text{AIE}}$ decreases, but large price perturbations distort production and would be likely be viewed unfavorably by various market stakeholders. The simplest guidance for choosing the size of the price perturbation is to choose a value of $c$ so that for the sample size of the experiment, the size of the price perturbation is noticeable to market participants but limited to a small percentage of the current market price (e.g. 1-2%). For markets where it is not possible for reputational or legal reasons to introduce any individual-level price variation, the components of $\tau^{*}_{\text{AIE}}$ can be estimated using price variation across products or time instead, at the cost of introducing additional structural assumptions on the environment.
In Theorem 2, we found that $\tau^{*}_{\text{AIE}}$ can be expressed in terms of certain price elasticities and the direct effect of treatment on net demand. Our first result using randomized price-shifters is that, using data collected under Design 2, we can produce a simple and consistent plug-in estimator for the indirect effect based on the functional form for $\tau^{*}_{\text{AIE}}$ derived in Theorem 2. Let $\bm{Y}$ be the $n$-length vector of observed outcomes, where $Y_{i}=Y_{i}(W_{i},P_{n}(\bm{W}),+U_{i}),$ $\bm{U}$ is the $n\times J$ matrix of price perturbations, and $\bm{Z}$ is the $n\times J$ matrix of observed net demand, where $Z_{i}=Z_{i}(W_{i},\bm{P}_{n}(\bm{W})+U_{i})$. The estimator is
@@
We provide a central limit theorem for our indirect effect estimator below. Its rate of convergence depends on the magnitude of the price perturbations $h_{n}$, and is always slower than the $\sqrt{n}$-rate obtained for the direct effect. In deriving this result, it is helpful to note that algorithmically $\hat{\tau}_{\text{AIE}}$ is the product of a rescaled instrumental variables estimator and a differences in mean estimator. We can then use standard results on the asymptotic behavior of IV estimators to guide our analysis.
@@
Suppose the Assumptions of Theorem 2 hold and that treatment and price-shifters are assigned according to Design 2. Then, the estimated average indirect effect can be written as:
@@ 4.1 Definitions of Conditional Treatment Effects Under Interference
is the expected effect of treating an individual with covariate value $x$ on everyone else’s outcomes in a system of $n$ individuals. These estimands are non-random quantities that depend both on the population distribution and on the market size. They are policy-relevant for settings where we expect to deploy a policy in a market of similar size and composition to the observed market. When interference is removed, the CADE is equal to the CATE, and the CAIE is zero.
@@
The next step, in Theorem 10, is to derive the population versions of the CADE and CAIE. These are purely population estimands, defined as the limit of $\bar{\tau}_{\text{CADE}}$ and $\bar{\tau}_{\text{CAIE}}$ as the market size grows to infinity. We will show below that these quantities are relevant to treatment targeting in a market that is larger but otherwise similar to the observed market. The population estimands are also useful for deriving estimators that are consistent for both the large-market and finite-market estimands. Theorem 10 shows that the population estimands have a simple and interpretable form and implies that estimators for one class of estimands are consistent for the other.
@@
The population estimands have a simple form that suggests estimation strategies for the estimands $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$. The limit of the CADE is the direct treatment effect on outcomes conditional on $x$, holding the equilibrium price fixed. The limit of the CAIE is the direct treatment effect on net demand conditional on $x$, multiplied by an elasticity correction that does not depend on $x$. The augmented randomized experiment from Design 2 can be used to estimate the elasticity corrections and the conditional average treatment effects required to estimate $\bar{\tau}_{\text{CADE}}(x)$ and $\bar{\tau}_{\text{CAIE}}(x)$.
@@ 4.2 Equilibrium-Stable Targeting
The above reveals that we can solve for the optimal price-stable policy without access to price elasticities. Thus, we can obtain a plug-in estimate for the optimal rule using CADE estimates derived from an RCT without price perturbations as in Design 1; recall that the CADE itself can be identified without price perturbations (Theorem 11). Furthermore, following Dantzig & Wald (1951), we can verify that its solution takes on a simple parametric form: The optimal price-stable policy is a thresholding rule that compares the CADE for the outcomes against a shadow cost of net demand effects $c\cdot\tau_{\text{CADE}}^{*,z}(x)$, where $c\in\mathbb{R}^{J}$ can be interpreted as a shadow-price vector.
@@
Under the conditions of Proposition 12, there exists a vector $c\in\mathbb{R}^{J}$ for which the asymptotic equilibrium-stable targeting problem (36) admits a solution $\nu(\cdot)$ with the following property: $\nu(x)=1$ whenever $\tau^{*}_{\text{CADE}}(x)>c^{\top}\,\tau_{\text{CADE}}^{*,z}(x)$ and $\nu(x)=0$ whenever $\tau^{*}_{\text{CADE}}(x)<c^{\top}\,\tau_{\text{CADE}}^{*,z}(x)$. Furthermore, if there exists $b\in[0,1]$ for which the policy
$\nu_{c,b}(x)=\begin{cases}1&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)>c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ b&\text{if}\ \ \tau^{*}_{\text{CADE}}(x)=c^{\top}\,\tau_{\text{CADE}}^{*,z}(x),\\ 0&\text{else,}\end{cases}$ (37)
satisfies the constraints in (36), then this policy is optimal.
We can estimate the optimal equilibrium-neutral rule by finding a rule that meets an empirical version of the conditions in Theorem 13, where $\tau^{*}_{\text{CADE}}(x)$ and $\tau_{\text{CADE}}^{*,z}(x)$ are replaced with appropriate estimators. With consistent estimators of the conditional average direct effects, such as the estimator proposed in Theorem 11, then the estimated equilibrium-neutral rule is consistent for the population equilibrium-neutral rule under some smoothness assumptions on the distribution of conditional average treatment effects. A description of the estimation procedure and a formal consistency result is provided in Appendix B.12.
@@ 5 Example: Cash-Transfer Experiments
The authors argue that this phenomenon is due to spillover effects, which largely arise from equilibrium effects in the market for perishable protein sources. Specifically, the authors provide evidence that eggs, which are a commonly consumed perishable protein source in the Philippines, have higher prices in treated compared to control villages where saturation of eligible households is high. This price effect is explained by the increase in demand for protein from eligible families and inelastic supply in remote villages. The higher prices in treated villages lead ineligible children to consume less protein, which explains a significant proportion of the negative effect of the cash transfer program on ineligible children. It thus appears that Assumption 2 is reasonable, in that the spillovers described in detail by the authors are restricted to those that occur through market prices.
@@ 5.1 Treatment Effects in a Village Economy
Under the assumption that markets in remote villages are isolated from one another, the village-level randomized experiment allowed Filmer et al. (2023) to recover treatment effects that include equilibrium effects. The total effect $\tau_{\text{TOT}}$ could be estimated using data from the cluster-randomized experiment by comparing the average outcomes of all children in treated compared to control villages. In this paper, in contrast, our focus is on randomized experiments where only a single market is observed, but randomization is possible at the household level. While our econometric setting is different from that of Filmer et al. (2023), we can use their dataset to conduct a calibrated evaluation of our approach. We first use moments from the cluster-randomized experiment in Filmer et al. (2023) to estimate a model of supply, demand, and children’s outcomes in a single remote village market. Then, we use this model to simulate a single-village, unit-randomized experiment, and assess the ability of our estimators of the average direct and indirect treatment effects to recover equilibrium effects from this simulated experiment.
@@
Given a treatment vector $\bm{W}$, the realized household-level outcomes are $Y_{h}(\bm{W})=Y_{h}(W_{h},P(\bm{W}))$ and household level net demand is $Z_{h}(\bm{W})=D_{h}(\bm{W},P(\bm{W}))-S(P(\bm{W}))$.
The 10 parameters in this model are estimated by matching 10 moments from a subset of the data in Filmer et al. (2023), restricted to households in villages above the 70th percentile of the remoteness index defined in the paper, where equilibrium effects were most meaningful. The 10 moments are:
@@
In Table 1, we show that both the direct and the indirect effect are estimable in finite samples using our proposed unit level augmented experiment. The size of the price-shifters in the simulated experiment is 0.15 Philippine pesos per egg, which is less than 2.5% of the market price. We report the results of a Monte Carlo simulation with 1,000 repetitions and a sample size of 2,000 households to evaluate the bias, variance, and coverage properties of the estimators and confidence intervals for $\hat{\tau}_{\text{ADE}}$ and $\hat{\tau}_{\text{AIE}}$. The table illustrates that the bias of the average direct effect and indirect effect estimator based on the augmented randomize experiment are low in finite samples. Furthermore, the coverage when $\hat{\sigma}_{D}$ and $\hat{\sigma}_{I}$ are used to construct confidence intervals in finite samples is slightly above the asymptotic confidence level of 95% for both population estimands. As expected from Theorem 3, intervals that target $\tau^{*}_{\text{ADE}}$ are conservative for $\tau_{\text{ADE}}$. For the AIE, in simulations the coverage for the sample estimand is also more conservative than for the population estimand. In this simple linear model, $\nabla_{p}y(1,p)=\nabla_{p}y(0,p)$, so $\Delta_{i}(W_{i},p)=0$ in (17), and confidence intervals for differences in means estimation that ignore price fluctuations are asymptotically valid. The augmented randomized experiment, however, is required to estimated the AIE in this model and to build asymptotically valid confidence intervals for the ADE in more general settings with additional heterogeneity and non-linearity.